# Andreas Schlaginhaufen SYCAMORE, EPFL

### v1

@@ D.7 Perturbation bounds
$\ell_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\leq\dfrac{1}{2\eta}\left\lVert[r^{\textsf{E}}]_{\mathcal{U}_{P}}-[\hat{r}]_{\mathcal{U}_{P}}\right\rVert_{2}^{2}\leq\dfrac{\hat{\varepsilon}}{\eta\sigma_{\mathcal{R}}\sin\left({\theta_{2}({P^{0}},{P^{1}})/2}\right)^{2}}.$ (95)
@@
$\ell_{P}\left({r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r})}\right)$ $\leq\dfrac{1}{2\eta}\left\lVert[r^{\textsf{E}}]_{\mathcal{U}_{P}}-[\hat{r}]_{\mathcal{U}_{P}}\right\rVert_{2}^{2}$ (98)
@@
Furthermore, the bound on the maximal principal angle follows from Proposition 3.4. ∎
Appendix H Proof of Theorem 1
@@
* The proof of Theorem 1 is inspired by [Syed and Schapire, 2007, Theorem 2]. However, in contrast to Syed and Schapire [2007], we consider the regularized problem with multiple experts, we use the suboptimality as the convergence metric, and we use a projected gradient descent update (instead of multiplicative weights). The proof hinges on Hoeffding’s inequality and a regret bound for online gradient descent, which are provided in Theorem H and 101 below.
@@
Step 2: When analyzing Algorithm 1 there are three sources of stochasticity. The first two are due to the randomness in the data sets $\mathcal{D}^{\textsf{E}}_{k}$ and $\mathcal{D}_{k,t}$, and the third is due to the randomness in the forward RL algorithm, $\mathsf{A}^{\varepsilon_{\text{opt}},\delta_{\text{opt}}}_{P^{k}}$, that upon a query with the reward $r_{t}$ outputs a policy $\pi_{k,t}$ such that with probability at least $1-\delta_{\text{opt}}$ it holds $\ell_{P^{k}}(r_{t},\mu^{\pi_{k,t}})\leq\varepsilon_{\text{opt}}$. Let’s denote the event that $\max_{r\in\mathcal{R}}\sum_{k=0}^{K-1}\langle r,\mu^{\textsf{E}}_{P^{k}}-\hat{\mu}_{\mathcal{D}^{\textsf{E}}_{k}}\rangle>\varepsilon_{\text{stat},\textsf{E}}$ by $\mathcal{E}_{\text{stat},\textsf{E}}$, the event that $\max_{r\in\mathcal{R}}\sum_{k=0}^{K-1}\langle r,\mu^{\pi_{k,t}}-\hat{\mu}_{\mathcal{D}_{k,t}}\rangle>\varepsilon_{\text{stat}}$ by $\mathcal{E}_{\text{stat},t}$, and the event that $\ell_{P^{k}}(r_{t},\mu^{\pi_{k,t}})>\varepsilon_{\text{opt}}$ by $\mathcal{E}_{\text{opt},k,t}$. Moreover, let us assume that $\mathcal{E}_{\text{stat},\textsf{E}}$ happens with probability at most $\delta_{\text{stat},\textsf{E}}$, $\mathcal{E}_{\text{stat},t}$ happens with probability at most $\delta_{\text{stat}}$, and $\mathcal{E}_{\text{opt},k,t}$ happens with probability at most $\delta_{\text{opt}}$. By union bound, the probability of the event
@@
Here, the inequalities $(i),(iv)$, and $(vi)$ follow from the concentration bound established in step 1. Moreover, inequality $(ii)$ holds since $\hat{r}\mapsto\max_{\mu\in\mathcal{M}_{P^{k}}}\langle\hat{r},\mu-\mu^{\textsf{E}}_{P^{k}}\rangle-\bar{h}(\mu)+\bar{h}(\mu^{\textsf{E}}_{P^{k}})$ is the pointwise maximum of affine functions and therefore convex. Furthermore, $(iii)$ follows from $\varepsilon_{\text{opt}}$-optimality of $\mu_{k,t}$, $(v)$ from Theorem 101, and $(vii)$ from concavity of the mapping $\mu_{k,t}\mapsto\langle r,\mu_{k,t}-\mu^{\textsf{E}}_{P^{k}}\rangle-\bar{h}(\mu_{k,t})$. Finally, $(viii)$ holds because all experts are optimal for the reward $r^{\textsf{E}}$. In conclusion, to ensure that with probability at least $1-\hat{\delta}$ it holds that $\ell_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\sum_{k=0}^{K-1}\ell_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\hat{\varepsilon}$ it suffices to choose $T=\frac{256K^{2}}{\hat{\varepsilon}^{2}}$, $\alpha=\frac{\hat{\varepsilon}}{16K^{2}}$, $N=\frac{128K\log\left({6|\mathcal{S}||\mathcal{A}|/\hat{\delta}}\right)}{\hat{\varepsilon}^{2}}$, $H=H_{t}=\frac{\log\left({16K/\hat{\varepsilon}}\right)}{\log(1/\gamma)}$, $N_{t}=\frac{128K\log\left({1536K^{2}|\mathcal{S}||\mathcal{A}|/(\hat{\delta}\hat{\varepsilon}^{2})}\right)}{\hat{\varepsilon}^{2}}$, $\delta_{\text{opt}}=\frac{\hat{\delta}\hat{\varepsilon}^{2}}{768K^{3}}$ , $\varepsilon_{\text{opt}}=\frac{\hat{\varepsilon}}{4K}$. ∎
Appendix I Experimental details
@@
(a)                         (b)                         (c)                         (d)
Figure 3: (a) shows the second principal angle between $P_{\beta}^{0}$ and $P_{\beta}^{1}$ for varying wind strength $\beta$. Furthermore, (b) shows the distance between $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ for a varying number of expert demonstrations $N^{\textsf{E}}$ and wind strength $\beta$. Moreover, (c) and (d) show the transferability to $P^{\text{South}}$ and $P^{\text{Shifted}}$ in terms of $\ell_{P^{\text{South}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{South}}}(\hat{r}))$ and $\ell_{P^{\text{Shifted}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{Shifted}}}(\hat{r}))$, respectively. The dots indicate the median and the shaded areas the 0.2 and 0.8 quantiles over the 10 independent realizations.

### v2

@@ D.7 Perturbation bounds
$\mathsf{SubOpt}_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\leq\dfrac{1}{2\eta}\left\lVert[r^{\textsf{E}}]_{\mathcal{U}_{P}}-[\hat{r}]_{\mathcal{U}_{P}}\right\rVert_{2}^{2}\leq\dfrac{\hat{\varepsilon}}{\eta\sigma_{\mathcal{R}}\sin\left({\theta_{2}({P^{0}},{P^{1}})/2}\right)^{2}}.$ (91)
@@
$\mathsf{SubOpt}_{P}\left({r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r})}\right)$ $\leq\dfrac{1}{2\eta}\left\lVert[r^{\textsf{E}}]_{\mathcal{U}_{P}}-[\hat{r}]_{\mathcal{U}_{P}}\right\rVert_{2}^{2}$ (94)
@@
Appendix H Estimating principal angles
Consider two full rank matrices $A,B\in\mathbb{R}^{n\times m}$ and let the columns of $U_{A},U_{B}\in\mathbb{R}^{n\times m}$ form an orthonormal basis of $\mathcal{V}=\operatorname{im}A$ and $\mathcal{W}=\operatorname{im}B$, respectively. Then, as discussed by [Ji-Guang, 1987] we have
$\sigma_{i}=\cos(\theta_{i}(\mathcal{V},\mathcal{W})),i=1,\ldots,m,$ (97)
where $1\geq\sigma_{1}\geq\ldots\geq\sigma_{m}\geq 0$ denote the singular values of $U_{A}^{\top}U_{B}$ sorted in decreasing order. Hence, given the transition matrices $P^{0},P^{1}$, we can compute the principle angles $\theta_{i}(P^{0},P^{1})$ by first computing orthonormal bases for the column spans of $E-\gamma P_{i},i=1,2$, and then computing the singular values as described above.
Now, suppose that $\hat{P}^{0},\hat{P}^{1}$ are empirical estimates of $P^{0},P^{1}$, then we have by [Ji-Guang, 1987, Theorem 3.1] the following perturbation result
$\max_{i}\left|{\sin(\theta_{i}(P^{0},P^{1}))-\sin(\theta_{i}(\hat{P}^{0},\hat{P}^{1}))}\right|$ $\leq\left\lVert\Pi_{\mathcal{U}_{P^{0}}}-\Pi_{\mathcal{U}_{\hat{P}^{0}}}\right\rVert+\left\lVert\Pi_{\mathcal{U}_{P^{1}}}-\Pi_{\mathcal{U}_{\hat{P}^{1}}}\right\rVert$ (98)
$\leq\gamma H_{\gamma}\sqrt{|\mathcal{S}|/|\mathcal{A}|}\left({\left\lVert P^{0}-\hat{P}^{0}\right\rVert+\left\lVert P^{1}-\hat{P}^{1}\right\rVert}\right),$ (99)
where the last inequality follows from Propositions D.8 and 3.4. Hence, we can estimate $\sin\theta_{i}(P^{0},P^{1})$ up to an error of $\mathcal{O}\left({\max\left\{{\left\lVert P^{0}-\hat{P}^{0}\right\rVert,\left\lVert P^{1}-\hat{P}^{1}\right\rVert}\right\}}\right)$.
Appendix I Proof of Theorem 1
@@
* The proof of Theorem 1 is inspired by [Syed and Schapire, 2007, Theorem 2]. However, in contrast to Syed and Schapire [2007], we consider the regularized problem with multiple experts, we use the suboptimality as the convergence metric, and we use a projected gradient descent update (instead of multiplicative weights). The proof hinges on Hoeffding’s inequality and a regret bound for online gradient descent, which are provided in Theorem I and 100 below.
@@
Step 2: When analyzing Algorithm 1 there are three sources of stochasticity. The first two are due to the randomness in the data sets $\mathcal{D}^{\textsf{E}}_{k}$ and $\mathcal{D}_{k,t}$, and the third is due to the randomness in the forward RL algorithm, $\mathsf{A}^{\varepsilon_{\text{opt}},\delta_{\text{opt}}}_{P^{k}}$, that upon a query with the reward $r_{t}$ outputs a policy $\pi_{k,t}$ such that with probability at least $1-\delta_{\text{opt}}$ it holds $\mathsf{SubOpt}_{P^{k}}(r_{t},\mu^{\pi_{k,t}})\leq\varepsilon_{\text{opt}}$. Let’s denote the event that $\max_{r\in\mathcal{R}}\sum_{k=0}^{K-1}\langle r,\mu^{\textsf{E}}_{P^{k}}-\hat{\mu}_{\mathcal{D}^{\textsf{E}}_{k}}\rangle>\varepsilon_{\text{stat},\textsf{E}}$ by $\mathcal{E}_{\text{stat},\textsf{E}}$, the event that $\max_{r\in\mathcal{R}}\sum_{k=0}^{K-1}\langle r,\mu^{\pi_{k,t}}-\hat{\mu}_{\mathcal{D}_{k,t}}\rangle>\varepsilon_{\text{stat}}$ by $\mathcal{E}_{\text{stat},t}$, and the event that $\mathsf{SubOpt}_{P^{k}}(r_{t},\mu^{\pi_{k,t}})>\varepsilon_{\text{opt}}$ by $\mathcal{E}_{\text{opt},k,t}$. Moreover, let us assume that $\mathcal{E}_{\text{stat},\textsf{E}}$ happens with probability at most $\delta_{\text{stat},\textsf{E}}$, $\mathcal{E}_{\text{stat},t}$ happens with probability at most $\delta_{\text{stat}}$, and $\mathcal{E}_{\text{opt},k,t}$ happens with probability at most $\delta_{\text{opt}}$. By union bound, the probability of the event
@@
Here, the inequalities $(i),(iv)$, and $(vi)$ follow from the concentration bound established in step 1. Moreover, inequality $(ii)$ holds since $\hat{r}\mapsto\max_{\mu\in\mathcal{M}_{P^{k}}}\langle\hat{r},\mu-\mu^{\textsf{E}}_{P^{k}}\rangle-\bar{h}(\mu)+\bar{h}(\mu^{\textsf{E}}_{P^{k}})$ is the pointwise maximum of affine functions and therefore convex. Furthermore, $(iii)$ follows from $\varepsilon_{\text{opt}}$-optimality of $\mu_{k,t}$, $(v)$ from Theorem 100, and $(vii)$ from concavity of the mapping $\mu_{k,t}\mapsto\langle r,\mu_{k,t}-\mu^{\textsf{E}}_{P^{k}}\rangle-\bar{h}(\mu_{k,t})$. Finally, $(viii)$ holds because all experts are optimal for the reward $r^{\textsf{E}}$. In conclusion, to ensure that with probability at least $1-\hat{\delta}$ it holds that $\mathsf{SubOpt}_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\sum_{k=0}^{K-1}\mathsf{SubOpt}_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\hat{\varepsilon}$ it suffices to choose $T=\frac{256K^{2}}{\hat{\varepsilon}^{2}}$, $\alpha=\frac{\hat{\varepsilon}}{16K^{2}}$, $N=\frac{128K\log\left({6|\mathcal{S}||\mathcal{A}|/\hat{\delta}}\right)}{\hat{\varepsilon}^{2}}$, $H=H_{t}=\frac{\log\left({16K/\hat{\varepsilon}}\right)}{\log(1/\gamma)}$, $N_{t}=\frac{128K\log\left({1536K^{2}|\mathcal{S}||\mathcal{A}|/(\hat{\delta}\hat{\varepsilon}^{2})}\right)}{\hat{\varepsilon}^{2}}$, $\delta_{\text{opt}}=\frac{\hat{\delta}\hat{\varepsilon}^{2}}{768K^{3}}$ , $\varepsilon_{\text{opt}}=\frac{\hat{\varepsilon}}{4K}$. ∎
Appendix J Suboptimal experts
In our problem formulation, we assumed that the $K$ experts are optimal with respect to $r^{\textsf{E}}$, i.e. $\mu^{\textsf{E}}_{P^{k}}=\mathsf{RL}_{P^{k}}(r^{\textsf{E}})$ for $k=0,\ldots,K-1$. This assumption can be weakened by requiring that
$\max_{r\in\mathcal{R}}\left|{J(r,\mu^{\textsf{E}}_{P^{k}})-J(r,\mathsf{RL}_{P^{k}}(r^{\textsf{E}}))}\right|\leq\varepsilon_{\text{mis}},$ (111)
where $\varepsilon_{\text{mis}}$ is some misspecification error. The transferability results in Theorem 3.4 and 3.4 still apply whenever we recover a reward $\hat{r}$ such that $\mathsf{SubOpt}_{P^{k}}(\hat{r},\mathsf{RL}_{P^{k}}(r^{\textsf{E}}))\leq\hat{\varepsilon}$. Moreover, with a straightforward modification of the proof of Theorem 1, it follows that with high probability Algorithm 1 recovers a reward $\hat{r}$ such that $\mathsf{SubOpt}_{P^{k}}(\hat{r},\mathsf{RL}_{P^{k}}(r^{\textsf{E}}))\leq\hat{\varepsilon}+2K\varepsilon_{\text{mis}}$. Hence, our end-to-end transferability guarantees apply with $\hat{\varepsilon}\leftarrow\hat{\varepsilon}+2K\varepsilon_{\text{mis}}$. However, $\varepsilon_{\text{mis}}$ cannot be further reduced by collecting more samples from the expert or MDP.
Appendix K Experimental details
@@
(a)                         (b)                         (c)                         (d)
Figure 4: (a) shows the second principal angle between $P_{\beta}^{0}$ and $P_{\beta}^{1}$ for varying wind strength $\beta$. Furthermore, (b) shows the distance between $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ for a varying number of expert demonstrations $N^{\textsf{E}}$ and wind strength $\beta$. Moreover, (c) and (d) show the transferability to $P^{\text{South}}$ and $P^{\text{Shifted}}$ in terms of $\mathsf{SubOpt}_{P^{\text{South}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{South}}}(\hat{r}))$ and $\mathsf{SubOpt}_{P^{\text{Shifted}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{Shifted}}}(\hat{r}))$, respectively. The dots indicate the median and the shaded areas the 0.2 and 0.8 quantiles over the 10 independent realizations.